OpenStudy (anonymous):

I completely forgot what an inequality is. can someone explain it to me please. :)

OpenStudy (studygurl14):

An inequality is a equation that uses the following signs instead of equal signs: \[<, >, \le, \ge\]

OpenStudy (studygurl14):

\(\large\sf\color{red}{For\ example: }\) \(\large 2<4\) \(\large x \ge 5\)

OpenStudy (anonymous):

Okay thank you so much for explaining it to me :D that helped a lot.

OpenStudy (studygurl14):

You're welcome. \(\large\ddot\smile\)